OpenStudy (anonymous):

Dividing Radical Expressions Simplify. 3/(4+4√5)

OpenStudy (anonymous):

multiply and divide by \[4-4\sqrt(5)\]

OpenStudy (anonymous):

How do I do that?

OpenStudy (jdoe0001):

do you know what a "conjugate" is?

OpenStudy (anonymous):

\[\frac{ 3 }{ 4+4\sqrt{5} }\times \frac{ 4-4\sqrt{5} }{ 4-4\sqrt{5} }\]

OpenStudy (anonymous):

\[\left( 4+4\sqrt{5} \right)\left( 4-4\sqrt{5} \right)=?\] use (a+b)(a-b)=a^2-b^2

OpenStudy (jdoe0001):

\(\bf \cfrac{3}{4+4\sqrt{5}}\cdot \cfrac{4-4\sqrt{5}}{4-4\sqrt{5}}\implies \cfrac{3(4-4\sqrt{5})}{(4+4\sqrt{5})(4-4\sqrt{5})} \\ \quad \\ recall\implies {\color{blue}{ (a-b)(a+b) = a^2-b^2}} \\ \quad \\ \cfrac{3(4-4\sqrt{5})}{(4+4\sqrt{5})(4-4\sqrt{5})}\implies \cfrac{3(4-4\sqrt{5})}{(4)^2-(4\sqrt{5})^2}\)